# OpenACE: An Open Benchmark for Evaluating Audio Coding Performance

### v1

@@ III-A Objective evaluation of audio quality vs bitrates
We leverage the intrusive VISQOL Audio [39] and POLQA [40] metrics for the objective evaluation. Given the 48 kHz sample rate requirement of the VISQOL algorithm, we select the full-band signals in the benchmark and resample them from 44.1 to 48 kHz using Kaiser window resampling. For objective metric computation, we truncate the audio to the first 10 seconds of each file. We evaluate the four coding algorithms at 16, 32, and 64 kbps.
Table III presents the objective scores for the various coding methods at the three different bitrates. The results show that at the low bitrate setting of 16 kbps, LC3+ shows the lowest performance, with LC3 performing slightly better. Notably, there is a significant gap of 0.4 between the two LC3-based codecs. EVS and Opus, on the other hand, achieve similar, higher results, with Opus having the highest score. Our results complement the previous LC3 assessment [38] that did not report the results obtained at 16 kbps. It is also fair to add that LC3 complexity encoding/decoding is significantly lower than Opus complexity. For example, in Hearing Aid music task, Opus has about twice as much complexity. This, however, results in degraded LC3 quality in lower bitrates.
@@ III-B Subjective evaluation of emotional speech at 16 kbps
Fourteen assessors conducted the MUSHRA test through an online platform [^2] https://senselabonline.com/ using the headphones. Tab. IV shows obtained quality aggregated over all tested codecs. Type III ANOVA test revealed that there is a statistically significant difference in emotion encoding quality (F value = 7.4, $p<0.001$). We thus confirm that audio coding degrades signal quality, as reported by the EMO-Codec benchmark [42]. Their authors found that both DSP and ML-based codecs do not encode emotional speech sufficiently, which is more challenging in lower (below 24 kbps) bitrates. Besides, the used IEMOCAP dataset [43] needs to be licensed, which is not required with the OpenACE data.

### v2

@@ III-A Objective evaluation of audio quality vs bitrates
We leverage the intrusive VISQOL Audio [39] and POLQA [40] metrics for the objective evaluation. NISQA [41] that supports full-band signals is also available in the project repository. Given the 48 kHz sample rate requirement of the VISQOL algorithm, we select the full-band signals in the benchmark and resample them from 44.1 to 48 kHz using Kaiser window resampling. For objective metric computation, we truncate the audio to the first 10 seconds of each file. We evaluate the four coding algorithms at 16, 32, and 64 kbps.
Table III presents the objective scores for the various coding methods at the three different bitrates. The results show that at the low bitrate setting of 16 kbps, LC3+ shows the lowest performance, with LC3 performing slightly better. Notably, there is a significant gap of 0.4 between the two LC3-based codecs. EVS and Opus, on the other hand, achieve similar, higher results, with Opus having the highest score. Our results complement the previous LC3 assessment [38] that did not report the results obtained at 16 kbps. It is also fair to add that LC3 complexity encoding/decoding is significantly lower than Opus complexity. For example, in Hearing Aid music task, Opus has about twice as much complexity.
@@ III-B Subjective evaluation of emotional speech at 16 kbps
Using the headphones, 36 listeners conducted the MUSHRA test through an online platform [^2] https://senselabonline.com/. Tab. IV shows obtained quality aggregated over all tested codecs. Type III ANOVA test revealed that there is a statistically significant difference in emotion encoding quality (F value = 8, $p<0.001$). We thus confirm that audio coding degrades signal quality, as reported by the EMO-Codec benchmark [43]. Their authors found that both DSP and ML-based codecs do not encode emotional speech sufficiently. Besides, the used IEMOCAP dataset [44] needs to be licensed, which is not required with the OpenACE data. The test data is available here [^3] https://huggingface.co/datasets/mcernak/EARS-EMO-OpenACE.